OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. m = -4, point (2, -5)

OpenStudy (anonymous):

plug in those given to the formula y-y1=m(x-x1) where (x1,y1)=2,-5) and m=-4, lets see what you get

OpenStudy (anonymous):

y-(-5)=-4x-2

OpenStudy (anonymous):

hmm theres a little slpped on the right hand side m(x-x1)

OpenStudy (anonymous):

never mind its just the bracket thing there lol..:D y-(-5)=-4(x-2) is it? y+5 isnt it?

OpenStudy (anonymous):

so my answer would be y-(-5)=-4(x-2)?

OpenStudy (anonymous):

thats also correct but you have to simplify them too.. like this one y-(-5)=y+5 and then multiply to simplify this -4(x-2)

OpenStudy (anonymous):

so then it would be y+5(-6x)?

OpenStudy (anonymous):

:(